OpenStudy (itsmichelle29):

What is the solution of log(3x + 4) 4096 = 4?

OpenStudy (anonymous):

4/3 would you like explanation

OpenStudy (itsmichelle29):

Nahhhh thanks

OpenStudy (campbell_st):

well raise both sides as powers of the base remember the log law for base \[\log_{a} (a^b) = b\] then 4096 = (3x + 4)^4 now you can solve for x

OpenStudy (itsmichelle29):

thanks @_greatmath7 @campbell_st

OpenStudy (anonymous):

write the equation (exponential) and then look... log(3x+4)4096=4⇒(3x+4)4=4096 8^4 = 4096 => 3x+4 = 8 (base must be positive) i got the answer 4/3?

OpenStudy (anonymous):

yeah

OpenStudy (itsmichelle29):

@campbell_st i would give u a medal but i cant to two people

OpenStudy (anonymous):

yw :)

OpenStudy (campbell_st):

@_greatmath7 one of the aims of open study is to help understanding, and try to avoid giving answers.

OpenStudy (anonymous):

oh I am sorry but I did also show how to solve so, if itsmichelle29 followed my steps she would have received the same answer

OpenStudy (anonymous):

@campbell_st